The following sample observations were randomly selected.

X 4 5 3 6 10

Y 4 6 5 7 7


Determine the regression equation.                                                        (6 marks)




1
Expert's answer
2021-11-09T16:31:00-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2\\
 \hline
 & 4 & 4 & 16 & 16 & 16\\
 & 5 & 6 & 30 & 25 & 36\\
 & 3 & 5 & 15 & 9 & 25\\
 & 6 & 7 & 42 & 36 & 49\\
 & 10 & 7 & 70 & 100 & 49\\
 Sum= & 28 & 29 & 173 & 186 & 175\\
\end{array}""\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{28}{5}=5.6"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{29}{5}=5.8"

"SS_{XX}=\sum_i(X_i-\bar{X})^2=29.2"

"SS_{YY}=\sum_i(Y_i-\bar{Y})^2=6.8"

"SS_{XY}=\sum_i(X_i-\bar{X})(Y_i-\bar{Y})=10.6"

"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{10.6}{29.2}=0.363014"

"n=\bar{Y}-m\bar{X}"

"=5.8-0.363014(5.6)"

"=3.767123"

Therefore, we find that the regression equation is:


"Y=3.767123+0.363014X"



Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}"

"=\dfrac{10.6}{\sqrt{29.2}\sqrt{6.8}}=0.752246"

We have strong positive correlation.

The regression equation is:


"Y=3.767123+0.363014X"